1) For Nominal Interest Rate expression select one of the following:

A. Interest period≥ Compounding period

B. Interest period = Compounding period

C. Interest period < Compounding period

D. Compounding period≥ Interest period.

2)If the nominal interest rate is 10% per year compounded weekly, then what is “z% per half year compounded weekly”?

A. 10%

B. 2%

C. 5%

D. 6%
Answer #1

1.

Option A is correct

Interest period≥ Compounding period

2.

APR = 10% compounded weekly

Weekly rate = 0.10/52 = 0.001923

Half yearly Rate = 0.001923(26) = 5%